OpenStudy (anonymous):

the difference between six times a number and 9 is equal to five times the sum of the number and 2. find the number step by step plz

OpenStudy (hoblos):

it can be written as: 6x - 9 = 5(x+2) => 6x -5x = 10+9 => x=19